### v1

# JunoBench: A Benchmark Dataset of Crashes in Python Machine Learning Jupyter Notebooks

Abstract.
Jupyter notebooks are widely used for machine learning (ML) prototyping. Yet few debugging tools are designed for ML code in notebooks, potentially due to the lack of benchmarks. We introduce JunoBench, the first benchmark dataset of real-world crashes in Python-based ML notebooks. JunoBench has 111 curated and reproducible crashes from public Kaggle notebooks, each paired with a verifiable fix, ranging over popular ML libraries, including TensorFlow/Keras, PyTorch, Scikit-learn, Pandas, and NumPy, as well as notebook-specific out-of-order execution issue. To support reproducibility and ease of use, JunoBench offers a unified execution environment where crashes and fixes can be reliably reproduced. By providing realistic crashes and their resolutions, JunoBench facilitates bug detection, localization, and repair tailored to the interactive and iterative nature of notebook-based ML development.
Benchmark, software bugs, machine learning, Jupyter notebooks
$c$s: Software and its engineering Software testing and debugging
1. Introduction
Machine learning (ML) and deep learning (DL) are increasingly used in modern software programs, with Jupyter notebooks emerging as a dominant development environment for ML prototyping (NASA high-end computing capability, 2025; Grotov et al., 2022). Notebooks offer a flexible, interactive interface that supports incremental coding, intermediate feedback, and out-of-order execution. These features make notebooks well suited for exploratory ML development, but also introduce unique debugging challenges (Grotov et al., 2022). Even worse, the use of complex ML libraries, such as TensorFlow and PyTorch, make ML prototyping extra error prone (Islam et al., 2019).
Existing tools offer limited support for identifying ML bugs, especially in the notebook environment (De Santana et al., 2024). One critical missing component is benchmarks. Benchmarks are crucial for advancing automatic debugging research by providing standardized datasets of real-world bugs for reproducible evaluation (v. Kistowski et al., 2015). While prior work (Morovati et al., 2023; Liang, 2022; Kim et al., 2021; Zhang et al., 2018; Wardat et al., 2021) focuses on ML bugs in Python scripts (.py files), they do not account for the unique semantics of Jupyter notebooks. Furthermore, existing ML bug benchmarks (e.g., (Morovati et al., 2023)) mix bug symptoms such as bad performance, incorrect functionality, and crashes. Crashes which cause code execution to terminate, are particularly valuable for benchmarking in the context of notebooks, as they are clearly observable from error outputs in .ipynb files.
To fill this gap, we present JunoBench, the first benchmark of crashes in real-world Python ML notebooks. Based on our previous empirical study (Wang et al., 2025c), JunoBench (available at Hugging Face (Wang et al., 2025a)) supports the evaluation of debugging tools for notebook-based ML development by providing:
  • 111 curated, reproducible real-world crashes from Python ML notebooks, each paired with a verifiable fix;
  • Broad, balanced coverage of major ML libraries (e.g., TensorFlow/Keras, PyTorch, Scikit-learn, Pandas, NumPy, Matplotlib, Seaborn) and Notebook-specific failures (out-of-order execution);
  • A unified execution environment for consistent reproduction and evaluation.
2. Methodology
This section outlines the methodology (see Figure 1), including data collection, benchmark construction, and how we ensure reproducibility and usability of the benchmark.
Figure 1. Overview of the benchmark construction process.:
2.1. Data collection
Data sources
The crashing notebooks in JunoBench were drawn from the notebook dataset in our previous empirical study of public Python ML notebook crashes (Wang et al., 2025c) where we studied 64,031 crashing notebooks and manually labeled 746 crashes with detailed annotations, including crash type, root cause, and the ML library used at the point of failure (referred to as the library cause).
The original notebook pool comprises notebooks collected from GitHub and Kaggle. This paper focuses specifically on Kaggle notebooks (2,869 notebooks containing 3,875 crashes, of which 356 are labeled) due to their greater support for reproducibility. For instance, Kaggle notebooks typically reference input datasets that are publicly accessible.
Crash filtering
We began with the 356 labeled crashes, and selected crashes that 1) occurred when using ML libraries (based on the library cause label), or 2) were due to notebook-specific out-of-order execution (the root cause label).
Then we excluded crashes whose root causes were not reproducible or fixable within the notebook context, such as 1) insufficient resource (hardware-dependent), 2) incompatible library versions (requiring additional complex environment setups), 3) bugs inside third-party libraries. Finally, 174 candidate crashes remained for the benchmark construction process.
Benchmark resampling
The filtered labeled crashes are unevenly distributed across ML libraries. To balance coverage, we resampled 500 additional crashes from the unlabeled Kaggle pool, selecting notebooks using at least one of the ML libraries used in the labeled set (i.e., TensorFlow/Keras, PyTorch, Scikit-learn, Pandas, NumPy, Matplotlib, Seaborn, Statsmodels, Torchvision, LightGBM, and OpenCV).
Due to the non-linear execution of notebooks, it is often difficult to trace the origin of variables in a target code cell and identify the actual ML libraries involved without manual inspection. Therefore, we implemented a tool to automatically estimate the library cause for each crash. The tool assigns scores to ML libraries based on their presence in the crashing code cell and its error traceback, giving higher weight to those appearing in the crashing line. The tool achieved an accuracy of 80% when evaluated against labeled data. It was used as a guide to prioritize which crashes to manually reproduce and fix first for benchmark inclusion.
2.2. Benchmark inclusion
Input dataset retrieval
To reproduce each crash, we first retrieved the necessary input datasets used by the notebooks. We used the KGTorrent (Quaranta et al., 2021) database and the Meta Kaggle (Risdal and Bozsolik, 2022) dataset to obtain dataset titles, downloading links, and licenses.
We then attempted to download the datasets using the downloading links. Since we focus on crashes instead of performance issues, the use of realistic or full training data is not relevant.Hence, for datasets over 100MB, we checked if downsampling preserved the crash behavior without altering code logic. For instance, a 2.5GB image dataset was reduced to 94MB by retaining just 4% of images per class while still reproducing the crash.
Manual reproducing and fixing procedure
In our previous study, the dataset was organized by crashes, with each entry representing a distinct crash rather than a notebook. Consequently, multiple crashes could originate from the same notebook. In constructing this benchmark, we adopted a different design in which each instance is represented as an independent notebook containing a single, isolated, reproducible crash. This ensures that each failure scenario is self-contained, simplifying evaluation and supporting consistent comparisons of automated debugging methods.
For each notebook, we keep the original notebook and create two modified versions. One is a reproduced version in which we apply only the minimal necessary code changes (e.g., paths or training settings). The other is a fixed version where we manually allocate and repair the root cause.
To reproduce each crash, we manually execute only a subset of code cells necessary to trigger the error. Due to non-linear notebook execution, we isolate each crash using one valid execution sequence, although alternatives may exist. This sequence is reflected in the cell execution counts, where higher numbers indicate later execution. In the fixed version, we preserve the same execution order for comparability. A crash is considered fixed when it no longer occurs and the notebook’s original intent is maintained.
Benchmark balancing across ML libraries
We began with 174 labeled crash candidates and successfully reproduced and fixed 74 of them. To balance ML library coverage, we continued with the additionally sampled 500 unlabeled ones. We first applied our previously introduced heuristic tool to automatically infer the library cause of each crash, based on which we prioritized our reproduction and fix efforts.To avoid overrepresentation of any single library, we monitored the distribution as we processed. Once a library became sufficiently represented, we halted further reproduction efforts for crashes associated with that library and focused on others. From this unlabeled set, we evaluated 93 additional crashes and successfully reproduced and fixed 37. Unreproducible crashes were primarily due to missing or inaccessible datasets.
In total, we curated 111 reproducible crashes and their verifiable fixes, covering both notebook-specific out-of-order execution issues, and balance the set of diverse ML libraries.
Crash categorization
For each included crash, we ensured consistent and accurate labeling to facilitate future downstream analysis and debugging tool development. For the 74 labeled crashes, we double-checked the labels, especially the root causes based on our manually verified fixes. We identified 7 cases with incorrect root cause labels. For example, one broadcast crash was originally labeled as data confusion, but it actually stemmed from an incorrect value being passed to the lag argument in the plot_acf API of the Statsmodels library, indicating the true root cause was API misuse. All mislabeled cases can be found in our GitHub repository (Wang et al., 2025b).
For the 37 unlabeled crashes, we manually labeled them following the approach in our empirical study (Wang et al., 2025c). Each crash was annotated with four labels: 1) crash type (e.g., tensor shape mismatch, unsupported broadcast), 2) root cause (e.g., API misuse, data confusion), 3) ML pipeline stage (where in the general ML pipeline the crash occurs; see 2(c)), and 4) library cause which was automatically inferred by our heuristic tool and manually validated in this step. Similar to our previous study, we applied grounded theory methods (Seaman, 1999), using first-cycle coding (labeling by one annotator) and second-cycle coding (review by another) (Saldana, 2015). Disagreements were resolved through discussion. In total, we changed 2 labels for ML pipeline stage and 3 for root cause after reviewing. We reached full agreement on crash type and library cause.
Input dataset license
Most datasets permit research use. In four cases where the original datasets (all numeric) are under a private or restrictive license, we replaced them with synthetic datasets by replicating the original dataset’s schema and generating random values. Each input dataset in JunoBench (Wang et al., 2025a) is accompanied by a metadata file that documents its title, source URL, license, and modifications applied (e.g., downsampling, synthetic substitution).
2.3. Benchmark usability
To improve usability, we provide a unified execution environment for all notebooks in JunoBench. We built a Docker image based on the official Kaggle notebook platform (Kaggle, 2025), extended it with additional dependencies required by the notebooks. This environment ensures all cases can be reproduced consistently with a one-time setup. The full dependency specification is provided in our benchmark repository (Wang et al., 2025a).
Furthermore, we offer an execution-level (Zhu et al., 2025) benchmark interface. It is a CLI tool that automatically runs the buggy or fixed version of each notebook in the reproduced cell execution order, and confirms whether the buggy version crashes as expected, while the fixed version executes successfully.
3. JunoBench
We analyze JunoBench in four dimensions: library cause, root cause, crash type, and ML pipeline stage.
(a) Library
(b) Root cause
(c) ML pipeline stage
Figure 2. Overview of library, root cause, and ML pipeline stage distribution in JunoBench. Each bar in (b) and (c) is segmented by libraries (a). “TF/K” stands for “tensorflow/keras”. “Minor Libs” include statsmodels (2), torchvision (1), and lightgbm (1).:
Library cause
JunoBench includes 111 notebook crashes with balanced coverage across popular ML libraries. It also covers notebook-specific issues. The library distribution is shown in 2(a). Specifically, JunoBench includes:
  • 15 crashes each related to: DL libraries TensorFlow/Keras and PyTorch, classical ML library Scikit-learn, and data processing libraries Pandas and NumPy;
  • 16 crashes covering a mix of ML libraries: 12 are visualization-related (6 from Matplotlib and 6 from Seaborn), and individual cases from Statsmodels, Torchvision, and LightGBM;
  • 20 notebook-specific (i.e., out-of-order execution) crashes.
This distribution highlights the benchmark’s diverse coverage of both challenges in using ML libraries and those arising from the flexible execution model of notebooks. It includes deep learning, classical ML, data processing, and visualization libraries, as well as issues such as out-of-order execution.
Root cause
The root causes of crashes covered by JunoBench are shown in 2(b). In addition to the 20 notebook-specific failures (i.e., out-of-order execution), such as re-running a code cell that deletes a column, resulting in an error when the column is referenced again, most of the remaining crashes arise from three categories. The most common is API misuse (45 cases), where developers invoke library functions with incorrect arguments or in unsupported ways. Data confusion follows (25), stemming from misunderstandings of data shape, type, or structure. Implementation errors (18) account for mistakes in code logic, such as incorrect variable usage or flawed algorithm design.
JunoBench also captures less common but notable real-world ML notebook issues. For example, ML model confusion (2 cases) occurs when loading a model that does not match the expected architecture defined in the code, or due to cross-framework confusion (e.g., using TensorFlow/Keras APIs on PyTorch models). Another rare case involves deprecated APIs removed in newer versions.
Crash type
Crashes in JunoBench span a variety of crash types. The most frequent is Invalid Argument (21), where function calls violate API constraints due to incorrect or missing parameters. Tensor-related errors such as Tensor Shape Mismatch (16) and Unsupported Broadcast (5) are also common, particularly when using TensorFlow/Keras and PyTorch where shape alignment is challenging. Data Value Violation (7) occurs when inputs contain unexpected or invalid values that break API or user assumptions. Another notable category is Variable Not Found (10), all caused by out-of-order cell execution, resulting in variables, defined in unexecuted code cells, becoming undefined.
These crash types capture how notebook developers frequently struggle with runtime dynamics and data integrity. ML-specific crash types, including tensor shape mismatch, data value violation, unsupported broadcast, reflect domain-specific challenges in ML notebook development.
ML pipeline stage
JunoBench captures crashes across key stages of a typical ML pipeline (2(c)). Most crashes occur during model training (31 cases), primarily involving TensorFlow/Keras and PyTorch, reflecting the complexity of using DL libraries. Data preparation follows closely (30), often failing from out-of-order execution, reflecting challenges of managing state in interactive notebook environments.
Model evaluation and prediction account for 21 crashes, often involving incorrect metric usage, or incompatible model inputs. Data visualization introduces 20 crashes, which not only involves libraries like Matplotlib and Seaborn but also reflects upstream data processing challenges in Pandas and NumPy. Finally, model construction is responsible for 9 crashes, typically caused by misconfigured architectures or invalid layer definitions in DL frameworks.
4. Potential research opportunities
JunoBench paves the way for advancing fault detection, localization, and repair in Python-based ML notebooks. It consists of 111 curated and reproducible real-world crashes Each crash includes buggy and fixed versions with necessary inputs, and can be reproduced in a unified environment. This supports analysis at both cell and line levels, and enables controlled and replicable tool development.
Crashes in Junobench are categorized by ML libraries, facilitating studies on how bugs vary across libraries or how well debugging techniques generalize across APIs. It also provides opportunities to develop library-specific debugging tools. Furthermore, unlike script-based benchmarks, JunoBench captures notebook execution metadata (e.g., cell execution order), enabling research on execution-aware analysis and tooling tailored to the notebook environment.
Evaluating large language models (LLMs) on JunoBench
With the increasing adoption of LLMs for software debugging, we evaluated two LLMs (i.e., llama3.3:70b and mistral-small3.1) as crash detectors in JunoBench at the cell level. Given all successfully executed code cells, each model is tasked with predicting whether a target cell will crash, boiling down to a binary classification problem. The target cell corresponds to the crashing cell in the reproduced buggy version (ground truth: TRUE) and the same cell in the fixed version (ground truth: FALSE).
JunoBench contains 111 buggy-fixed notebook pairs, but some inputs exceed the models’ token limits (11 pairs for llama3.3:70b and 3 for mistral-small3.1), resulting in 200 and 216 total predictions, respectively. Performance is reported as an aggregate over five runs using a Majority vote strategy. llama3.3:70b correctly identified crashes in 115 out of 200 cases (buggy: 58/100, fixed: 57/100), while mistral-small3.1 achieved 131 correct predictions out of 216 (buggy: 65/108, fixed: 66/108). These low scores reflect the difficulty of the crash prediction task on our benchmark, suggesting that it presents a meaningful challenge for current LLM-based debugging approaches. Details about the experiments can be found in our GitHub repository (Wang et al., 2025b).
5. Related Work
Several studies have introduced bug benchmarks in Python-based ML applications. Morovati et al. (Morovati et al., 2023) propose Defect4ML, a benchmark of 100 reproducible bugs in ML systems using TensorFlow/Keras. The bugs were collected from GitHub, Stack Overflow, and prior studies (Zhang et al., 2018; Wardat et al., 2021; Nikanjam et al., 2021; Islam et al., 2020). Liang et al. (Liang, 2022) present gDefects4DL, a curated set of 64 DL bugs in TensorFlow/Keras and PyTorch projects. However, their download links for bugs and replication environments are no longer active, limiting their usability. Kim et al. (Kim et al., 2021) introduce Denchmark, a large-scale dataset comprising 4,577 bug reports from DL projects. While it offers extensive coverage, it lacks reproduction-ready artifacts, making it unsuitable for studies requiring replication or tool evaluation.
Several ML studies have released bug datasets as byproducts. Zhang et al. (Zhang et al., 2018) compiled 151 reproduced TensorFlow bugs in their empirical study. Wu et al. (Wu et al., 2021) introduce Tensfa for tensor shape mismatches in TensorFlow/Keras applications. A key limitation of the datasets by Zhang et al. (Zhang et al., 2018) and Wu et al. (Wu et al., 2021) is their reliance on outdated Python and TensorFlow versions without precise dependency details, reducing their relevance and reproducibility for modern ML development (Jahangirova et al., 2024).
All of the above studies target ML applications written in Python scripts (.py file) instead of Jupyter notebooks (.ipynb). While several notebook datasets exist, they focus on general notebook usage (Grotov et al., 2022; Quaranta et al., 2021), reproducibility (Wang et al., 2021), or quality analysis (Pimentel et al., 2019; Quaranta et al., 2022), and do not include bug benchmarks. De Santana et al. (De Santana et al., 2024) conducted an empirical study of bugs in Jupyter notebook projects. Their dataset lacks reproducible artifacts such as actual buggy and fixed notebook versions, environments, or test inputs.
In contrast, JunoBench targets crashes in ML notebooks, providing reproducible buggy and fixed notebook pairs and input data. To our best knowledge, it is the first benchmark dataset of ML notebook crashes. Furthermore, while prior benchmarks (Morovati et al., 2023; Liang, 2022; Zhang et al., 2018; Wu et al., 2021) center on TensorFlow/Keras and PyTorch, JunoBench spans a wider range of ML libraries, capturing popular libraries used in real-world ML development. Moreover, while many existing benchmarks (Morovati et al., 2023; Liang, 2022; Wardat et al., 2021) require an individual environment setup per case, we provide a unified execution environment to ensure that all cases can be reproduced consistently with a one-time setup.
6. Conclusion
In this paper, we present JunoBench, the first benchmark dataset of real-world crashes in Python-based ML notebooks. JunoBench consists of 111 curated and reproducible crashes from Kaggle notebooks, each paired with a verifiable fix. JunoBench covers a balanced range of popular ML libraries (i.e., TensorFlow/Keras, PyTorch, Scikit-learn, Pandas, and Numpy), and notebook-specific out-of-order execution crash type. To promote ease of use, JunoBench provides a unified execution environment that enables consistent reproduction of all crashes and verification of corresponding fixes.
In the future, we plan to continuously expand and maintain JunoBench by adding crashes and updating library versions to reflect ongoing changes in ML development. We also aim to integrate JunoBench with debugging tools, including static analyzers and LLMs, to foster research into notebook-aware debugging solutions.
Acknowledgements.
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and the Software Center Project 61.